# Zebrafix: Mitigating Memory-Centric Side-Channel Leakage via Interleaving

### v1

@@ Abstract
Constant-time code has become the de-facto standard for secure cryptographic implementations. However, some memory-based leakage classes such as ciphertext side-channels, silent stores, and data memory-dependent prefetching remain unaddressed. In the context of ciphertext side-channel mitigations, the practicality of interleaving data with counter values remains to be explored.
To close this gap, we define design choices and requirements to leverage interleaving for a generic ciphertext side-channel mitigation. Based on these results, we implement Zebrafix, a compiler-based tool to ensure freshness of memory stores. We evaluate Zebrafix and find that interleaving can perform much better than other ciphertext side-channel mitigations, at the cost of a high practical complexity. We further observe that ciphertext side-channels, silent stores and data memory-dependent prefetching belong to a broader attack category: memory-centric side-channels. Under this unified view, we discuss to what extent ciphertext side-channel mitigations can be adapted to prevent all three memory-centric side-channel attacks via interleaving.
@@ 1 Introduction
In the second part of this work, we show that ciphertext side-channels are actually part of a broader category of memory-centric side-channel attacks: Data that has been written to memory once leaks secret information through subsequent accesses to the unprotected data. By generalizing the leakages to memory-centric instead of fixing each leakage in isolation, future mitigations can benefit from existing knowledge and proven effective strategies for developing a mitigation. One illustrative example from the category of memory-centric side-channels is the upcoming microarchitectural optimization silent stores [24], which discards stores that would not result in a change to the memory contents. Consequently, this breaks a vital assumption of constant-time code, namely that the behavior of stores is independent of the value written. Another related example is data memory-dependent prefetching (DMP) [37, 21], where the CPU prefetches data depending on (potential) pointers that it detected in recently accessed memory.
We discuss similarities and differences between these attacks, and analyze the suitability of existing analysis and mitigation tools. For example, a performant compiler level mitigation against silent store leakage is presented in cio [19]. The addition of a memory write with a provably distinct value between two consecutive memory writes effectively prevents silencing a store. The cio approach does not cover ciphertext side-channel leakage.
On the side of ciphertext side-channel analysis tools, CipherH [15] aims to find leaky memory writes by reasoning whether two consecutive memory writes of the same secret variable to the same address can hold different values. We observe that this approach is a limited analysis scope in terms of ciphertext side-channel leakage and would result in false negative results for code that is protected with a cio-style silent store mitigation approach. There are currently no dedicated DMP leakage mitigations; yet, a Cipherfix-style masking approach is applicable to protect secret data at the cost of introducing high overheads. In light of this, we point out how existing ciphertext side-channel defenses can be generalized to mitigate silent store leakage. However, adapting interleaving to mitigate the entire class of attacks, including DMP, is non-trivial and invites future research focusing on more holistic mitigations.
1.1 Contribution
@@
  • We propose strategies for augmenting interleaving as a mitigation against other memory-centric side-channel leakages.
@@
Outline. We provide background on the different leakage channels in Section 2 and establish interleaving as a ciphertext side-channel mitigation in Section 3. We continue with a comparison of different mitigation implementation layers in Section 4. Then, in Section 5, we present our Zebrafix implementation, which we evaluate in Section 6. We explain the interconnections between ciphertext side-channel leakages and silent store leakages in Section 7, and DMP leakages in Section 8. We discuss potential enhancements and general technical limitations in Section 9 and set our work in context with related work in Section 10. We conclude with Section 11.
2 Leakage Preliminaries
@@
2.1 Constant-Time Swap
@@
2.2 Ciphertext Side-Channels
@@
The authors of [25] present two attacks that are based on the missing freshness of ciphertexts, dubbed dictionary attack and collision attack, whereby both rely on monitoring repeated write accesses to the same memory address. The dictionary attack enables the attacker to map a set of ciphertexts to a set of known plaintexts to learn information about secret data. In the stronger attack scenario, the collision attack, the attacker knows that there are only two different ciphertexts at a given address. From this, the attacker can infer whether the data being written is the same as the data already present at that address by learning an equality predicate of the data. When the attacker monitors the respective architectural state after multiple memory writes to certain addresses, they can infer whether the plaintext changed or not after each write access: The ciphertext at that address changes if and only if the plaintext is different. A graphical representation of ciphertext side-channel leakage through collision attacks is given in Figure 2. Ciphertext side-channel mitigations are further investigated in Section 3.
Figure 2: Ciphertext side-channel leakage in $\mathtt{ct}$-$\mathtt{swap}$ based on [41, p. 3]. Ciphertexts are depicted in different shades of the memory address block. Although an attacker can only see ciphertexts of the values in arrays $\mathtt{a}$ and $\mathtt{b}$, a ciphertext change leaks the value of the secret decision bit $\mathtt{s}$ during the $\mathtt{ct}$-$\mathtt{swap}$.
2.3 Silent Stores
@@
An attacker needs to be able to observe the microarchitectural state of system components (e.g., cache state, execution time or store queue pressure) in order to derive whether a store has been silenced or not. Silent stores can either lead to similar leakage as the ciphertext side-channel by observing whether data at a certain address changes or not, or it can lead to leakage by observing whether an attacker-controlled value and a secret value are equal and lead to the suppression of a silent store.
The leakage from this microarchitectural optimization can be circumvented by ensuring that no memory write to a specific address contains a value that is already present at that address, e.g., by adding a write with new data in between or applying software-based probabilistic encryption to the data before writing it to memory [19]. The size of the intermediate write in this scenario depends on the width of the store operation.
@@
2.4 Data Memory-Dependent Prefetchers
3 Preventing Ciphertext Side-Channels
To counter ciphertext side-channel leakage, masking has been thoroughly explored as a binary level mitigation, while interleaving has received little attention yet. In the following, we describe interleaving as a ciphertext side-channel mitigation in depth and discuss associated design decisions.
3.1 Threat Model
@@
3.2 Interleaving as Leakage Defense
While there are plans to add hardware-based read protections in AMD SEV-SNP for upcoming CPU generations [4], the currently available processors need a software approach for protection. There are three software techniques that can be employed to prevent ciphertext side-channel leakage [25, 42], each taking a different route at circumventing ciphertext determinism: First, to prevent repeating plaintexts at the same address, the reuse of memory locations can be limited. Second, plaintexts can be masked with random values to introduce software-based nondeterminism. Masking adds a random bit string before each store, which needs to be subtracted on loads. This raises certain issues, like the quality of the random number generator, which strongly correlates with the performance overhead. Both, rotating addresses of data and masking, were found to incur considerable performance overhead in practical implementations [25, 42].
@@ 3.2 Interleaving as Leakage Defense
In line with previous work, data types to be protected include primitive integers, and arrays or structs containing those types. As noted in [42], pointers do not need protection: The memory location of variables can be assumed to be known, and secret-dependent memory accesses are eliminated by the constant-time code. Typically, the protected data resides on the stack, the heap or in global variables.
@@
Common memory encryption schemes use AES, which has a 128-bit block size. Thus, each plaintext block has 16 bytes. The amount of bytes reserved for the counter is a security parameter, as it determines the minimum number of memory writes for which no collisions are possible. Given that the native integer size of most current systems is 64-bit, it is reasonable to divide each 16-byte block evenly into two 8-byte chunks. This guarantees $2^{64}$ writes of a 16-byte block without collisions. Increasing the counter size yields an even larger security margin, but would lead to splitting 8-byte numbers, which would increase implementation complexity and overhead considerably. This issue also applies when the counter size is decreased, as the opened space is smaller than 8 bytes. We thus conclude that a counter size of 8 bytes is optimal in terms of security and implementation complexity. An example for interleaving with 8-byte counters is shown in Figure 5.
@@ 3.3 Realization Considerations
Cipherfix takes the first approach: The authors modify the application source code to call a special dummy function, passing each secret as a parameter. They then employ dynamic taint tracking to find all instructions accessing that variable. Using dynamic methods for secrecy tracking has the main drawback that it is very susceptible to insufficient coverage. Missing a code path during analysis eventually leads to bugs in the protected application, where an instruction accesses a protected memory location under a wrong assumption. Static taint analysis tools may avoid this issue. Yet, static analyses often need to overapproximate results and can still be incomplete in the tracking of exact data types, leading to error-proneness in the instrumentation. Any secrecy tracking that leads to hybrid (instrumented and non-instrumented) usage of global variables endangers functional correctness of the program, as detailed in Section 9.2.
To avoid this hybrid usage, we take the second approach in Zebrafix, i.e., comprehensive function annotation. We protect all variables and computations in the call tree underlying the outmost caller function of the cryptographic routines, reducing the taint tracking problem to a simpler and broader call graph analysis that leads to protection of all called functions and all allocated heap memory. Function cloning can help to simplify the addition of interleaving only where needed, but would still have the problem of the necessary synchronization of global variables between instrumented and non-instrumented functions. Though instrumenting all functions in the call graph clearly overapproximates the protection scope (comparable to heavy overtainting), it simplifies the implementation and greatly reduces the error-proneness. While the intent of precise information flow tracking is reduction of the runtime performance overhead, we found that Zebrafix performs quite well, despite the overapproximation.
4 Defense Implementation Layers
As previous research indicated, ciphertext side-channel defenses can be realized at different implementation layers: The ad-hoc patches to WolfSSL [47] and the Linux kernel [25] were made to the source code, while Cipherfix [42] directly modified the compiled binaries through a binary rewriting framework.
In the following, we provide guidance on a suitable implementation layer for an automated mitigation. We accumulate knowledge gained from existing side-channel defenses, explore possible solutions and discuss practical limitations. For compiler-based approaches, we generally use terminology from the LLVM project [23], which has long become a standard tool for side-channel defense implementations. However, our conclusions also apply to similar compiler infrastructures.
4.1 Binary Rewriting
@@
Finally, precise variable tracking is mandatory, as pointer and type information is usually not available. Missing a particular code path may lead to underapproximation, which in turn results in spurious runtime crashes due to unexpected data layout. In summary, while binary rewriting was found to work in principle, it was deemed too unstable for production use [42].
4.2 Compiler: Machine IR/Back End
@@
In the back end, there are different stages of lowering the architecture-independent IR to MIR. In order to work with the back end code that already contains all the memory writes present in the compiled binary, the mitigation passes must operate after the register allocation is completed. Therefore, the instrumentation must either perform the same register allocation steps as the binary rewriting, or use an option to ignore certain registers at the cost of a global performance impact. Instrumentation must also take care of preserving architectural state like the status flags register, which can be changed by arithmetic operations such as mask application or counter increments. Similar to binary rewriting, tracking of variables may be necessary, if the MIR does not offer detailed pointer and type information from earlier compilation steps. We conclude that MIR is generally the better choice over binary rewriting in terms of stability and performance, if there are no dependencies that should be instrumented as well.
4.3 Compiler: LLVM IR/Middle End
@@
Summary. In order to benefit from eased adjustments to memory layouts, instrumentation needs to be on middle end level. The advantages of back end level instrumentation only become apparent if the instrumentation happens after the register allocation pass to be able to protect all memory writes. Yet, at the back end stage, it is already sufficiently complicated to adjust the memory layout for interleaving, making it comparable to binary instrumentation. We conclude that hardening ciphertext side-channels at the IR layer promises the best final performance; and combined with a binary level verification, an appropriate security assurance is provided.
5 Zebrafix Implementation
Figure 6: Zebrafix toolchain. The developer marks sensitive workloads via function annotations. Then, the LLVM middle end passes add the main interleaving capabilities to the code. In the LLVM back end, the register allocation is adapted to prefer spilling to vector registers over spilling to the stack.
@@ 5 Zebrafix Implementation
As a case study for evaluating the described interleaving approach, we implement a proof-of-concept, dubbed Zebrafix. With that, we showcase the versatility of interleaving as a side-channel defense. Zebrafix is implemented as a set of LLVM 16111LLVM 16 is at the time of writing the newest LLVM version supported by SVF. out-of-tree passes with SVF [35] pointer analysis, with 2600 LOC in total. An overview of the Zebrafix toolchain can be found in Figure 6. The developer who wants to protect a certain function against ciphertext side-channel leakage marks it with a $\mathtt{clang}$ function attribute. The attribute is automatically propagated through the call tree, leading to instrumentation of all called child functions as well. We merge all bitcode files that are generated for whole-program analysis and then instrument the merged bitcode file. Zebrafix supports the block size 16 byte with data chunks up to size $16/2=8$ byte. An example of data chunks smaller than 8 byte can be found in Figure 7.
5.1 Compiler Level Transformations
@@
We base our mitigation implementation mainly on LLVM IR passes, after all IR optimization has run. This allows us to work on optimized LLVM IR, but before register allocation. To keep the approach modular and to ease combination with other leakage mitigations, we adjust the data layout via IR passes, instead of manipulating the heap memory allocators via a runtime library. Thereby, we avoid expensive heuristics and also support allocations known at compile time, including global variables and stack variables.
Implementing the mitigation on LLVM IR level succeeds with the current standard configuration of LLVM to produce interleaved memory writes when the compilation process is otherwise adjusted correctly: For a working transformation with Zebrafix, vector registers need to be available on the target, and compile steps prior to the instrumentation passes must avoid using them. As the interleaved counter and the data must be written simultaneously per 16-byte block, vector writes are inserted. For each primitive type, we provide a crafted struct that contains the suitable data type, some dummy elements and the 8-byte counter. Adding new primitive types that have a suitable length would break compatibility with existing LLVM versions, including the analysis and optimization infrastructure [^2] https://llvm.org/docs/ExtendingLLVM.html. Working with existing primitive types in structs combined with inserting vector writes allows us to circumvent the insertion of new types so that we can also take advantage of the highly optimized LLVM infrastructure. Moreover, with the flexibility of using structs to fill data into memory, we can comprehensibly adjust to other leakage scenarios stemming from other current and upcoming microarchitectural optimizations (see Section 7 and Section 8).
@@
In addition to the LLVM IR passes, we adjust the register allocator for x86-64 targets to decrease the number of stack spills of secret data during function calls or because of high register pressure inside of functions. For that, we integrate a register allocator enhancement that spills to vector registers instead of the stack, proposed by Matthias Braun [^3] We adapted https://discourse.llvm.org/t/rfc-spill2reg-selectively-replace-spills-to-stack-with-spills-to-vector-registers/59630/15.. If there are not enough vector registers available for saving values, we print a warning so that the developer can check whether the write actually contains secret-dependent data or just irrelevant data like a pointer that is written to the stack. The check is further enhanced with a binary level verification tool.
5.2 Adjusting Compiler Intrinsics

### v2

@@ Abstract
Constant-time code has become the de-facto standard for secure cryptographic implementations. However, some memory-based leakage classes such as ciphertext side-channels and silent stores remain unaddressed. Prior work proposed three different methods for ciphertext side-channel mitigation, for which one, the practicality of interleaving data with counter values, remains to be explored. To close this gap, we define design choices and requirements to leverage interleaving for a generic ciphertext side-channel mitigation. Based on these results, we implement Zebrafix, a compiler-based tool to ensure freshness of memory stores. We evaluate Zebrafix and find that interleaving can perform much better than other ciphertext side-channel mitigations, at the cost of a high practical complexity. We further observe that ciphertext side-channels and silent stores belong to a broader attack category: memory-centric side-channels. Under this unified view, we show that interleaving-based ciphertext side-channel mitigations can be used to prevent silent stores as well.
@@ I Introduction
In the second part of this work, we show that ciphertext side-channels are actually part of a broader category of memory-centric side-channel attacks: Data that has been written to memory once leaks secret information through subsequent accesses to the unprotected data. By generalizing the leakages to memory-centric instead of fixing each leakage in isolation, future mitigations can benefit from existing knowledge and proven effective strategies for developing a mitigation. One illustrative example from the category of memory-centric side-channels is the upcoming microarchitectural optimization silent stores [20], which discards stores that would not result in a change to the memory contents. Consequently, this breaks a vital assumption of constant-time code, namely that the behavior of stores is independent of the value written. Another related example is data memory-dependent prefetching (DMP) [32, 17], where the CPU prefetches data depending on (potential) pointers detected in recently accessed memory.
We discuss similarities and differences between ciphertext side-channels and silent stores, and analyze the suitability of existing analysis and mitigation tools. For example, a performant compiler-level mitigation against silent store leakage is presented in cio [15]. The addition of a memory write with a provably distinct value between two consecutive memory writes effectively prevents silencing a store. The cio approach does not cover ciphertext side-channel leakage.
On the side of ciphertext side-channel analysis tools, CipherH [12] aims to find leaky memory writes by reasoning whether two consecutive memory writes of the same secret variable to the same address can hold different values. We observe that this approach has a limited analysis scope in terms of ciphertext side-channel leakage and would result in false negative results for code that is protected with a cio-style silent store mitigation approach. There are currently no dedicated DMP leakage mitigations; yet, a Cipherfix-style masking approach is applicable to protect secret data at the cost of introducing high overheads. In light of this, we point out how existing ciphertext side-channel defenses can be generalized to mitigate silent store leakage. However, adapting interleaving to mitigate the entire class of attacks, including DMP, is non-trivial and invites future research focusing on more holistic mitigations.
I-A Contribution
@@
  • We discuss strategies for augmenting interleaving as a mitigation against other memory-centric side-channel leakages.
@@
Outline. We provide background on the different leakage channels in Section II and establish interleaving as a ciphertext side-channel mitigation in Section III. We continue with a comparison of different mitigation implementation layers in Section IV. Then, in Section V, we present our Zebrafix implementation, which we evaluate in Section VI. We explain the interconnections between ciphertext side-channel leakages and silent store leakages in Section VII. We discuss potential enhancements and general technical limitations in Section VIII and set our work in context with related work in Section IX. We conclude with Section X. An additional discussion of data memory-dependent prefetching in relation to interleaving can be found in Appendix A.
II Leakage Preliminaries
@@
II-A Constant-Time Swap
@@
II-B Ciphertext Side-Channels
@@
The authors of [21] present two attacks that are based on the missing freshness of ciphertexts, dubbed dictionary attack and collision attack, whereby both rely on monitoring repeated write accesses to the same memory address. The dictionary attack enables the attacker to map a set of ciphertexts to a set of known plaintexts to learn information about secret data. In the stronger attack scenario, the collision attack, the attacker knows that there are only two different ciphertexts at a given address. From this, the attacker can infer whether the data being written is the same as the data already present at that address by learning an equality predicate of the data. When the attacker monitors the respective architectural state after multiple memory writes to certain addresses, they can infer whether the plaintext changed or not after each write access: The ciphertext at that address changes if and only if the plaintext is different. A graphical representation of ciphertext side-channel leakage through collision attacks is given in Figure 2. Ciphertext side-channel mitigations are further investigated in Section III.
Figure 2: Ciphertext side-channel leakage in $\mathtt{ct}$-$\mathtt{swap}$ based on [36, p. 3]. Ciphertexts are depicted in different shades of the memory address block. Although an attacker can only see ciphertexts of the values in arrays $\mathtt{a}$ and $\mathtt{b}$, a ciphertext change leaks the value of the secret decision bit $\mathtt{s}$ during the $\mathtt{ct}$-$\mathtt{swap}$.
II-C Silent Stores
@@
An attacker has to be able to observe the microarchitectural state of system components (e.g., cache state, execution time or store queue pressure) to derive whether a store has been silenced or not. Silent stores can either lead to similar leakage as the ciphertext side-channel by observing whether data at a certain address changes or not, or it can lead to leakage by observing whether an attacker-controlled value and a secret value are equal and lead to the suppression of a silent store.
The leakage from this microarchitectural optimization can be circumvented by ensuring that no memory write to a specific address contains a value that is already present at that address, e.g., by adding a write with new data in between or applying software-based probabilistic encryption to the data before writing it to memory [15].
@@
III Preventing Ciphertext Side-Channels
To counter ciphertext side-channel leakage, masking has been thoroughly explored as a binary-level mitigation, while interleaving has received little attention yet. In the following, we describe interleaving as a ciphertext side-channel mitigation in depth and discuss associated design decisions.
III-A Threat Model
@@
III-B Interleaving as Leakage Defense
While there are plans to add hardware-based read protections in AMD SEV-SNP for upcoming CPU generations [4], the currently available processors need a software approach for protection. There are three software techniques that can be employed to prevent ciphertext side-channel leakage [21, 37], each taking a different route at circumventing ciphertext determinism: First, to prevent repeating plaintexts at the same address, the reuse of memory locations can be limited. Second, plaintexts can be masked with random values to introduce software-based nondeterminism. Masking adds a random bit string before stores, which needs to be subtracted on loads. This raises certain issues, like the quality of the random number generator, which strongly correlates with the performance overhead. Both, rotating addresses of data and masking, were found to incur considerable performance overhead in practical implementations [21, 37].
@@ III-B Interleaving as Leakage Defense
In line with previous work, protected data types include primitive integers, and arrays or structs containing those types. As noted in [37], pointers do not need protection: The memory location of variables can be known, and secret-dependent memory accesses are eliminated by the constant-time code. Typically, the protected data resides on the stack, the heap or in global variables.
@@
Common memory encryption schemes use AES, which has a 128-bit block size. Thus, each plaintext block has 16 bytes. The amount of bytes reserved for the counter is a security parameter, as it determines the minimum number of memory writes for which no collisions are possible. Given that the native integer size of most current systems is 64-bit, it is reasonable to divide each 16-byte block evenly into two 8-byte chunks. This guarantees $2^{64}$ writes of a 16-byte block without collisions.
Increasing the counter size yields an even larger security margin, but would lead to splitting 8-byte numbers, which would increase implementation complexity and overhead considerably. This issue also applies when the counter size is decreased, as the opened space is smaller than 8 bytes. We thus conclude that a counter size of 8 bytes is optimal in terms of security and implementation complexity. An example for interleaving with 8-byte counters is shown in Figure 4.
@@ III-C Realization Considerations
Cipherfix takes the first approach: The authors modify the application source code to call a special dummy function, passing each secret as a parameter. They then employ dynamic taint tracking to find all instructions accessing that variable. Using dynamic methods for secrecy tracking has the main drawback that it is very susceptible to insufficient coverage. Missing a code path during analysis eventually leads to bugs in the protected application, where an instruction accesses a protected memory location under a wrong assumption. Static taint analysis tools may avoid this issue. Yet, static analyses often need to overapproximate results and can still be incomplete in the tracking of exact data types, leading to error-proneness in the instrumentation. Any secrecy tracking that leads to hybrid (instrumented and non-instrumented) usage of global variables endangers functional correctness of the program, as detailed in Section VIII.
To avoid this hybrid usage, we take the second approach in Zebrafix, i.e., comprehensive function annotation. We protect all variables and computations in the call tree underlying the outmost caller function of the cryptographic routines, reducing the taint tracking problem to a simpler and broader call graph analysis that leads to protection of all called functions and all allocated heap memory. Though instrumenting all functions in the call graph clearly overapproximates the protection scope (comparable to heavy overtainting), it simplifies the implementation and greatly reduces the error-proneness. While the intent of precise information flow tracking is reduction of the runtime performance overhead, we found that Zebrafix performs quite well, despite the overapproximation.
IV Defense Implementation Layers
As previous research indicated, ciphertext side-channel defenses can be realized at different implementation layers: The ad-hoc patches to WolfSSL [42] and the Linux kernel [21] were made to the source code, while Cipherfix [37] directly modified the compiled binaries through a binary rewriting framework. In the following, we provide guidance on a suitable implementation layer for an automated mitigation. We accumulate knowledge gained from existing side-channel defenses, explore possible solutions and discuss practical limitations. For compiler-based approaches, we generally use terminology from the LLVM project [19], which has long become a standard tool for side-channel defense implementations. However, our conclusions also apply to similar compiler infrastructures.
IV-A Binary Rewriting
@@
Finally, precise variable tracking is mandatory, as pointer and type information is not available. Missing a particular code path may lead to underapproximation, which in turn results in spurious runtime crashes due to unexpected data layout. In summary, while binary rewriting works in principle, it was deemed too unstable for production use [37].
IV-B Compiler: Machine IR/Back End
@@
In the back end, there are different stages of lowering the architecture-independent IR to MIR. In order to work with the back end code that already contains all the memory writes present in the compiled binary, the mitigation passes must operate after the register allocation is completed. Therefore, the instrumentation must either perform the same steps to free registers as the binary rewriting, or use an option to ignore certain registers at the cost of a global performance impact. Instrumentation must also take care of preserving architectural state like the status flags register, which can be changed by arithmetic operations such as mask application or counter increments. Similar to binary rewriting, tracking of variables may be necessary, if the MIR does not offer detailed pointer and type information from earlier compilation steps. We conclude that MIR is generally the better choice over binary rewriting in terms of stability and performance, if there are no dependencies that should be instrumented as well.
IV-C Compiler: LLVM IR/Middle End
@@
Summary. In order to benefit from eased adjustments to memory layouts, instrumentation needs to be on middle end level. The advantages of back end level instrumentation only become apparent if the instrumentation happens after the register allocation to be able to protect all memory writes. Yet, at the back end stage, it is already sufficiently complicated to adjust the memory layout for interleaving, making it comparable to binary instrumentation. We conclude that hardening ciphertext side-channels at the IR layer promises the best final performance; and combined with a binary-level verification, an appropriate security assurance is provided.
V Zebrafix Implementation
Figure 5: Zebrafix toolchain. The developer marks sensitive workloads via function annotations. Then, the middle end passes add the main interleaving capabilities to the code. In the back end, the register allocation is adapted to prefer spilling to vector registers over spilling to the stack. With memory traces, the binary can be checked for leaky writes.
@@ V Zebrafix Implementation
As a case study for evaluating the described interleaving approach, we implement a proof-of-concept, dubbed Zebrafix. With that, we showcase the versatility of interleaving as a side-channel defense. Zebrafix is implemented as a set of LLVM 16111LLVM 16 is at the time of writing the newest LLVM version supported by SVF. out-of-tree passes with SVF [30] pointer analysis, with $2,600$ LOC in total. An overview of the Zebrafix toolchain can be found in Figure 5. The developer who wants to protect a certain function against ciphertext side-channel leakage marks it with a $\mathtt{clang}$ function attribute. The attribute is automatically propagated through the call tree, leading to instrumentation of all called child functions as well. Indirect calls and unsupported components like inline assembly are printed for further inspection by the developer during the call tree instrumentation which is based on the LLVM call graph. We merge all bitcode files that are generated for whole-program analysis and then instrument the merged bitcode file. A discussion about the interaction of instrumented and non-instrumented (e.g., dynamically linked) code can be found in Section VIII. Zebrafix supports the block size 16 byte with data chunks up to size $16/2=8$ byte. An example of data chunks smaller than 8 byte can be found in Figure 6.
V-A Compiler-Level Transformations
@@
We base our mitigation implementation mainly on LLVM IR passes, after all IR optimization has run. This allows us to work on optimized LLVM IR, but before register allocation. To keep the approach modular and to ease combination with other mitigations, we adjust the data layout via IR passes, instead of manipulating allocators and memory accesses via a runtime library.
Implementing the mitigation on LLVM IR level succeeds with the current standard configuration of LLVM to produce interleaved memory writes when the compilation process is otherwise adjusted correctly: For a working transformation with Zebrafix, vector registers need to be available on the target, and compile steps prior to the instrumentation passes must avoid using them. As the interleaved counter and the data must be written simultaneously per 16-byte block, vector writes are inserted. For each primitive type, we provide a crafted struct that contains the suitable data type, some dummy elements and the 8-byte counter. Adding own primitive types that have a suitable length would break compatibility with existing LLVM versions, including the analysis and optimization infrastructure [^2] https://llvm.org/docs/ExtendingLLVM.html#adding-a-new-type. Working with existing primitive types in structs combined with inserting vector writes allows us to circumvent the insertion of own types so that we can also take advantage of the highly optimized LLVM infrastructure. Moreover, with the flexibility of using structs to fill data into memory, we can comprehensibly adjust to other leakage scenarios stemming from other current and upcoming microarchitectural optimizations (see Section VII and Appendix A).
@@
In addition to the LLVM IR passes, we adjust the register allocator for x86-64 targets to decrease the number of stack spills of potentially secret data during function calls or because of high register pressure inside of functions. For that, we integrate a register allocator enhancement that spills to vector registers instead of the stack, proposed by Matthias Braun [^3] We adapted https://discourse.llvm.org/t/rfc-spill2reg-selectively-replace-spills-to-stack-with-spills-to-vector-registers/59630/15.. If there are not enough vector registers available for saving values, we print a warning so that the developer can check whether the write actually contains secret-dependent data or just irrelevant data like a pointer that is written to the stack. The check is further enhanced with a binary-level verification tool. For the sake of maintainability and extensibility of Zebrafix, we chose not to implement interleaved stack spilling, which would have required deep changes to the stack allocation.
V-B Adjusting Compiler Intrinsics